# The HITRAN2024 methane update

### v2

@@ 2.7.2 HITRAN update in the 6770–7570 cm-1 region
The new PLDL list [malarich_dual_2021] mostly replaces the 12CH4 transition wavenumbers and some lower energy levels of HITRAN2020 while keeping the old intensities. More than 22,000 transitions were updated using their data, and about 300 new lines were added. The improved residuals between the measurement and updated database are shown in Fig. 10. Similarly to the DLR/NIST/USTC measurements, the new pressure-induced parameters were used as described in Section 3.
@@ 2.8.1 IAO & LIPhy CRDS measurements
An extended assignment and analysis of 12CH4 lines in the 7606–7919 cm-1 region was recently reported by Nikitin et al. [nikitin_first_2024]. The experimental methane spectra in this range recorded with CRDS technique in Grenoble [mondelain_2011] at 80 K and 296 K correspond to the high-energy part of the WKLMC empirical line list [campargue_2013]. Some lines in this range belong to the upper edge of the Icosad band system previously considered in [rey_2015, rey_2016], but most of the other transitions have been included in HITRAN2020 without assignments.
Quantum assignments for 1382 transitions pointing to 33 vibrational sub-levels of the 12CH4 Triacontad system were obtained using a combined approach based on variational ab initio predictions [rey_2013, rey_new_2018] with a subsequent empirical optimization of the effective spectroscopic models derived from the potential and dipole moment energy surfaces as described in [tyuterev_2022, rey_2022]. They belong to three weak Triacontad bands, $6\nu_{4}$, $\nu_{2}$$5\nu_{4}$, $\nu_{1}$$4\nu_{4}$, falling in the methane transparency window near 1.28 µm. The positions of these vibrational bands are in good agreement with the values obtained from the ab initio potential energy surface [nikitin_var_2018, nikitin_pes_2025]. The root-mean-square fit deviation for the line positions of newly assigned transitions was 0.006 cm-1. The intensities of the line list for these bands span the range $10^{-25}$ to $9\times 10^{-30}$ cm/molecule. The full statistical results of the analyses are given in Table 1 of [nikitin_first_2024] and an overview plot of the 12CH4 experimental line lists near 1.28 µm is shown in Figure 4 of [nikitin_first_2024].
@@ 2.8.2 HITRAN update in the 7600–7920 cm-1 region
The line list provided by Nikitin et al. [nikitin_first_2024] completes the 12CH4 main band in the lower part of the triacontad. More than 400 new lines were added, and more than 50 were updated on their positions, intensities and lower state energy levels. Another major benefit of the update in that region comes with the addition of rotational assignments, which were previously very scarce and are now mostly complete for these bands. With $E_{\mathrm{Lower}}$, this allows for modeling in a wider range of temperatures and gives better estimation of line-shape parameters.
@@ 2.9.1 LIPhy Kitt peak FTS and CRDS measurements 10802–13922 cm-1
Figure 11: Update of the methane line list above 10800 cm-1. The HITRAN2020 list (orange circles) was completed by new FTS and CRDS measurements (blue and green circles, respectively). The TheoReTS variational list [rey_new_2018] (grey circles) is plotted as background below 13500 cm-1.
In Ref. [campargue_high_2023], empirical line lists at room temperature were retrieved from a long-path Kitt Peak FTS spectrum between 10802 and 13922 cm-1 (about 12800 lines) and from a high-sensitivity CRDS spectrum in the 13060–13300 cm-1 interval (about 2650 lines). Above 10800 cm-1, the previous HITRAN2020 dataset was limited to the 10923–11502 cm-1 interval where measurements obtained by intracavity laser absorption spectroscopy (ICLAS) were used as source (see Brown et al. [brown_methane_2013]). The FTS results of Ref. [campargue_high_2023] were found to be in good agreement with the HITRAN2020 dataset.
At these wavenumbers, the methane spectrum is especially highly congested even at room temperature, and, according to Ref. [campargue_high_2023], part of the absorption shows up as a “quasi-continuum”, which should be added to the absorption lines to fully account for the methane absorption in the region. The “quasi-continuum” represents about one third of the total absorption in the 11000–11400 cm-1 region. Note that the intensity sum of the lines and of the continuum was found to be in good agreement with the intensity predictions of the TheoReTS list [rey_new_2018]. In this context, one can use the weak lines from the HITEMP database [10.3847/1538-4365/ab7a1a] (which is largely based on TheoReTS, especially in this region).
@@ 2.9.2 HITRAN update in the 10800–14000 cm-1 region
This new list from Campargue et al. [campargue_high_2023] completes the update below 10923 cm-1 and above 11502 cm-1: regions that were previously empty (see Fig. 11), with an additional 5000 FTS and 1800 CRDS new transitions, and excluding the quasi-continuum, which contributes to the total absorption and will need to be accounted for in the future. These lines are new to HITRAN. As a consequence of being completely empirical, the new lines were attributed to the main isotopologue by default. This issue will have to be addressed in a future HITRAN update.
@@ 2.10 Overview
Line list Isotopologue
Transition wavenumbers
Intensities Region covered (cm-1)
Nikitin et al., 2025 [nikitin_improved_2024] 12CH4 1880 1610 900 – 1050
@@
ExoMol [yurchenko_exomol_2024] 12CH4 240 140 2503 – 4177
MeCaSDa [RICHARD2024109127] 12CH4 70 30 2786 – 3910
@@
Sung et al. [sung_new_2024] 13CH4 4710 4710 4000 – 4700
Nikitin et al., 2013 [nikitin_preliminary_2013] 12CH3D 1530 1510 4001 – 4554
Rodina et al., 2021 [rodina_improved_2021] 12CH4 4930 4230 4100 – 4190
@@
Ben Fathallah et al., 2024 [benfathallah_ch3d_2024] 12CH3D 10630 5850 6098 – 6530
Malarich et al., 2021 [malarich_dual_2021] 12CH4 22320 270 6770 – 7570
Nikitin et al., 2024 [nikitin_first_2024] 12CH4 480 480 7629 – 7919
Campargue et al., 2023 [campargue_high_2023] 12CH4 6780 6780 10803 – 13923
@@ 3.1 USTC, NIST & DLR measurements
For transitions in the $2\nu_{3}$ band of methane, HITRAN2020 only covers the parameters of Voigt profile (VP), i.e., pressure broadening ($\gamma_{0}$) and pressure shifting ($\delta_{0}$). Yet, to achieve greater precision in atmospheric methane simulations, more intricate line shapes and precise parameters are needed. In this regard, we employed the Speed-Dependent Voigt profile (SDVP) corrected for first-order Line-Mixing (LM), abbreviated as LMSDVP, as our base model to analyze the CRDS data. SDVP is a simplification of the pCqSDHC(HT) profile discussed in Ngo et al.[NGO2013JQSRT], which is more accurate than the VP owing to the inclusion of the speed-dependent effect. The involvement of numerous fitting parameters in the HT profile, compounded by the intricate nature of methane spectra, can result in closely numerically correlated spectral line parameters. This can impede the convergence of fitting solutions when analyzing the actual spectra or cause the derived physical parameters to be less physical. Therefore, we did not use the HT profile. Indeed, the SDVP is adequate for most endeavors of spectral analysis and simulations, but additional treatment is needed due to the high degeneracy of methane states. Thus, we introduced the first-order Rosenkranz parameter, $Y_{n}$, to account for line mixing effects [Rosenkranz1975]. From a computational perspective, $Y_{n}$ exhibits a strong correlation with line intensity $S$. Consequently, an independent measurement of line intensity is crucial for acquiring dependable line-mixing parameters. Together with the intensity measurements by NIST and DLR, a global-fit procedure was developed based on the LMSDV model. We also converted the measured nitrogen-broadening coefficients to air-broadening coefficients using the formula $\gamma_{0}^{\text{air}}=0.79\gamma_{0}^{\mathrm{N}_{2}}+0.21\gamma_{0}^{\mathrm{O}_{2}}$, and utilizing polynomial parameters for $\gamma_{0}^{\mathrm{O}_{2}}$ as provided by Lyulin et al. [Lyulin2009JQSRT]. The nitrogen pressure shifting coefficients were similarly scaled as $\delta_{0}^{\text{air}}=0.79\delta_{0}^{\mathrm{N}_{2}}+0.21\delta_{0}^{\mathrm{O}_{2}}$, again utilizing parameters for $\delta_{0}^{\mathrm{O}_{2}}$ as provided by Lyulin et al. [Lyulin2009JQSRT]. Values of $Y_{n}^{\mathrm{N\textsubscript{2}}}$ for the measured CH4–N2 system are converted for CH4–air through multiplication by 0.985, as shown in Tran et al. for the $2\nu_{3}$ band [Tran2010]. The speed-dependent Voigt parameters were incorporated in HITRAN2024 wherever available.
@@
Broadening parameters ($\gamma_{0}$) were gathered from all post-year 2000 sources [antony_n2_2008, bertin_co2_2024, clark_difference_2004, devi_self-_2015, devi_spectral_2016, devi_multispectrum_2018, es-sebbar_intensities_2014, es-sebbar_linestrengths_2021, farji_airinduced_2021, gabard_line_2010, gharib-nezhad_h2induced_2019, ghysels_spectroscopy_2011, ghysels_temperature_2014, grigoriev_estimation_2001, lyulin_measurements_2014, ma_temperature_2016, manne_determination_2017, mondelain_line_2005, mondelain_measurement_2007, pine_speeddependent_2019, pine_speeddependent_2000, pine_multispectrum_2003, richard_self_2023] that we could find, and separated by collision partners, transition symmetries, and ranking index. Unlike for other rotors, only the ranking index, $\alpha$, lets us differentiate between lines of the same manifold and symmetry, which simply relates to the energy level instead of a motion. Because of this, the parameters do not exhibit any clear pattern with $m$ for each index (see Fig. 14). Thus, separation was only used to isolate the most intense lines, which are found in $\alpha=1$.
@@
The different line profiles used in these studies impact the value of the retrieved coefficients by a small but non-negligible amount. For consistency with the parameters already present in HITRAN, we chose to focus on the Voigt profile. Separating each profile would greatly reduce the amount of data per fit; therefore, we scaled the respective non-Voigt parameters by a constant value to match the Voigt broadening. The constant scaling was calculated by taking the ratio of Voigt to non-Voigt broadening for each transition and averaging the result. This approach could only be performed in cases where transitions were fit with both non-Voigt and Voigt profiles —thus not all studies could be included. To further increase the amount of available data, the CH4–N2 pressure broadening parameters were combined with those of –air, and scaled by a factor of 0.988 taken from Pine & Gabard [pine_speeddependent_2000]. This approach was previously used in other studies [mondelain_measurement_2007] and agrees very well with other measurements (see Fig. 15). Finally, thanks to the weak branch dependence, broadening parameters for the $P$, $Q$, and $R$ branches were combined. Ideally, each branch should be separate; however, the broadening measurements in methane are too uncertain for the difference to be reliably quantified.
@@
With its different symmetry, CH3D must be treated differently, and benefits from a more rigorous separation based on quantum numbers. The semi-empirical approach developed by Dudaryonok et al. [dudaryonok_semiempirical_2018] seems to produce very good results in the $\nu_{6}$ band. Ignoring the wavenumber dependence, we applied their method for the broadening and temperature-dependence parameters of every new line added in Sec. 2. The wavenumber dependence of the pressure shift was reintroduced, summing the semi-empirical value to the empirical formula in HITRAN2012 [brown_methane_2013] of $2\times 10^{-6}\times\Delta\nu$, limited to negative values. A similar approach was used where the extremely low or high values of HITRAN2020 were replaced along with the new lines, which only amount to about 70.
@@ 4 Conclusion
The parameters of more than 80,000 lines of 12CH4, 13CH4, and 12CH3D have been updated for HITRAN20204 using the new line lists provided. Transitions wavenumbers, intensities, and lower-state energy levels were taken directly from these lists to replace the values from HITRAN2020. Laboratory spectra in the pentad, octad and tetradecad regions were used by the HITRAN team to validate and choose between the overlapping data sets, while other regions relied on the validations from the other groups. We observed a great drop in the residuals of the octad and tetradecad regions spectra particularly thanks to the atm.161 list and the new data from Reed et al.[reed_multilaboratory_2025] and Yan et al.. The effects on the pentad region were limited, but still showed improvements, especially at high er energies. 45,000 lines of the various isotopologues have been added, including some lower-intensity CH3D lines that should have been in the previous update but were missing, and increasing the region covered from 12,000 to 14,000 cm-1 with experimental data. In regions that were previously empirically determined, this work allowed us to reassign some of the intensities to the correct isotopologues. Assignments of many lines have been completed between the octad and triacontad regions.
The broadening parameters of these newly added lines were replaced using Padé-approximant functions fit from multiple recent studies for 12CH4 and 13CH4, unless reliable experimental data were provided, such as those generated at the USTC. The functions were obtained from data sets separated based on their rotational symmetry and ranking index. Currently, only one function per symmetry was used, regardless of the ranking index. The pressure-induced shifts were calculated using the same formula from HITRAN2012 [brown_methane_2013]. Semi-empirical functions from Dudaryonok et al. [dudaryonok_semiempirical_2018] were used for 12CH3D. The pressure shifts were scaled using the HITRAN2012 formula to account for the vibrational dependence, while the broadening parameters were taken as is. The broadening parameters from the $\nu_{3}$ band, were replaced similarly using the same Padé approximants, however, experimental values were used when available.
@@
The relational structure of the HITRAN database [Hill2016] and functionality of HAPI [Kochanov2016] allow one to add non-Voigt line shape parameters. For instance, in this edition, we have added the speed-dependent Voigt parameters and first-order line mixing for air-broadening in the tetradecad region. It will be important to have a consistent set of parameters in all the bands as was done for CO2 [Hashemi2020], N2O, and CO [Hashemi2021] in HITRAN2020. Of course, it is much harder for methane as there are strong symmetry and vibrational dependencies in comparison with linear molecules. Moreover, it will be important to introduce the full line mixing parameters, where possible. The accuracy of first-order line mixing deteriorates quickly at pressures approaching $1$ atm and above. Dense spectral regions such as $Q$ branches are especially affected and need full line mixing, possibly in combination with speed-dependent effects [ciurylo_speeddependent_2000, bertin_co2_2024], to be properly measured. Unlike some other parameters, these require complete knowledge of the assignments to be implemented; therefore, more theoretical work will have to be done in preparation. Currently, data from Tran et al. is available for CH4–air [Tran2006, Tran2010] and other collision partners [tran_model_2006-1, Tran2022], and will be considered in future updates.
In addition, in order to support the studies of planetary and brown dwarf atmospheres, it is important to provide parameters associated with broadening by “planetary” gases, including H2O [Tan2019], H2, He and CO2 [Tan2022]. Unlike many other gases for which these “planetary” broadening parameters are available, only broadening by water vapor is available for methane. In this work, we have used the scaling factors recommended in Ref. [Tan2019] to assign broadening by water vapor values to all new lines of methane. We have collected existing information on broadening by hydrogen, helium, and carbon dioxide (including Refs. [bertin_co2_2024, Tran2022, Varanasi1971, Varanasi1974, Varanasi1990, Fox1989, Gabard2013, es-sebbar_intensities_2014, gharib-nezhad_h2induced_2019, Vispoel2019, Sung2020, Yousefi2021, Hosokawa2025, Clment2025], but it is still a lot of work to process these data, which often do not agree with each other within stated uncertainties.

### v3

@@ 2.7.2 HITRAN update in the 6770–7570 cm-1 region
The new PLDL list [60] mostly replaces the 12CH4 transition wavenumbers and some lower energy levels of HITRAN2020 while keeping the old intensities. More than 22,000 transitions were updated using their data, and about 300 new lines were added. The improved residuals between the measurement and updated database are shown in Fig. 10. Similarly to the DLR/NIST/USTC measurements, the new pressure-induced parameters were used as described in Section 3.
@@ 2.8.1 IAO & LIPhy CRDS measurements
An extended assignment and analysis of 12CH4 lines in the 7606–7919 cm-1 region was recently reported by Nikitin et al. [72]. The experimental methane spectra in this range recorded with CRDS technique in Grenoble [64] at 80 K and 296 K correspond to the high-energy part of the WKLMC empirical line list [13]. Some lines in this range belong to the upper edge of the Icosad band system previously considered in [88, 85], but most of the other transitions have been included in HITRAN2020 without assignments.
Quantum assignments for 1382 transitions pointing to 33 vibrational sub-levels of the 12CH4 Triacontad system were obtained using a combined approach based on variational ab initio predictions [86, 84] with a subsequent empirical optimization of the effective spectroscopic models derived from the potential and dipole moment energy surfaces as described in [114, 90]. They belong to three weak Triacontad bands, $6\nu_{4}$, $\nu_{2}$$5\nu_{4}$, $\nu_{1}$$4\nu_{4}$, falling in the methane transparency window near 1.28 µm. The positions of these vibrational bands are in good agreement with the values obtained from the ab initio potential energy surface [70, 71]. The root-mean-square fit deviation for the line positions of newly assigned transitions was 0.006 cm-1. The intensities of the line list for these bands span the range $10^{-25}$ to $9\times 10^{-30}$ cm/molecule. The full statistical results of the analyses are given in Table 1 of [72] and an overview plot of the 12CH4 experimental line lists near 1.28 µm is shown in Figure 4 of [72].
@@ 2.8.2 HITRAN update in the 7600–7920 cm-1 region
The line list provided by Nikitin et al. [72] completes the 12CH4 main band in the lower part of the triacontad. More than 400 new lines were added, and more than 50 were updated on their positions, intensities and lower state energy levels. Another major benefit of the update in that region comes with the addition of rotational assignments, which were previously very scarce and are now mostly complete for these bands. With $E_{\mathrm{Lower}}$, this allows for modeling in a wider range of temperatures and gives better estimation of line-shape parameters.
@@ 2.9.1 LIPhy Kitt peak FTS and CRDS measurements 10802–13922 cm-1
Figure 11: Update of the methane line list above 10800 cm-1. The HITRAN2020 list (orange circles) was completed by new FTS and CRDS measurements (blue and green circles, respectively). The TheoReTS variational list [84] (grey circles) is plotted as background below 13500 cm-1.
In Ref. [12], empirical line lists at room temperature were retrieved from a long-path Kitt Peak FTS spectrum between 10802 and 13922 cm-1 (about 12800 lines) and from a high-sensitivity CRDS spectrum in the 13060–13300 cm-1 interval (about 2650 lines). Above 10800 cm-1, the previous HITRAN2020 dataset was limited to the 10923–11502 cm-1 interval where measurements obtained by intracavity laser absorption spectroscopy (ICLAS) were used as source (see Brown et al. [11]). The FTS results of Ref. [12] were found to be in good agreement with the HITRAN2020 dataset.
At these wavenumbers, the methane spectrum is especially highly congested even at room temperature, and, according to Ref. [12], part of the absorption shows up as a “quasi-continuum”, which should be added to the absorption lines to fully account for the methane absorption in the region. The “quasi-continuum” represents about one third of the total absorption in the 11000–11400 cm-1 region. Note that the intensity sum of the lines and of the continuum was found to be in good agreement with the intensity predictions of the TheoReTS list [84]. In this context, one can use the weak lines from the HITEMP database [42] (which is largely based on TheoReTS, especially in this region).
@@ 2.9.2 HITRAN update in the 10800–14000 cm-1 region
This new list from Campargue et al. [12] completes the update below 10923 cm-1 and above 11502 cm-1: regions that were previously empty (see Fig. 11), with an additional 5000 FTS and 1800 CRDS new transitions, and excluding the quasi-continuum, which contributes to the total absorption and will need to be accounted for in the future. These lines are new to HITRAN. As a consequence of being completely empirical, the new lines were attributed to the main isotopologue by default. This issue will have to be addressed in a future HITRAN update.
@@ 2.10 Overview
Line list Isotopologue Transition wavenumbers Intensities Region covered (cm-1)
Nikitin et al., 2025 [75] 12CH4 1880 1610 900 – 1050
@@
ExoMol [124] 12CH4 240 140 2503 – 4177
MeCaSDa [93] 12CH4 70 30 2786 – 3910
@@
Sung et al. [102] 13CH4 4710 4710 4000 – 4700
Nikitin et al., 2013 [76] 12CH3D 1530 1510 4001 – 4554
Rodina et al., 2021 [95] 12CH4 4930 4230 4100 – 4190
@@
Ben Fathallah et al., 2024 [5] 12CH3D 10630 5850 6098 – 6530
Malarich et al., 2021 [60] 12CH4 22320 270 6770 – 7570
Nikitin et al., 2024 [72] 12CH4 480 480 7629 – 7919
Campargue et al., 2023 [12] 12CH4 6780 6780 10803 – 13923
@@ 3.1 USTC, NIST & DLR measurements
For transitions in the $2\nu_{3}$ band of methane, HITRAN2020 only covers the parameters of Voigt profile (VP), i.e., pressure broadening ($\gamma_{0}$) and pressure shifting ($\delta_{0}$). Yet, to achieve greater precision in atmospheric methane simulations, more intricate line shapes and precise parameters are needed. In this regard, we employed the Speed-Dependent Voigt profile (SDVP) corrected for first-order Line-Mixing (LM), abbreviated as LMSDVP, as our base model to analyze the CRDS data. SDVP is a simplification of the pCqSDHC(HT) profile discussed in Ngo et al.[67], which is more accurate than the VP owing to the inclusion of the speed-dependent effect. The involvement of numerous fitting parameters in the HT profile, compounded by the intricate nature of methane spectra, can result in closely numerically correlated spectral line parameters. This can impede the convergence of fitting solutions when analyzing the actual spectra or cause the derived physical parameters to be less physical. Therefore, we did not use the HT profile. Indeed, the SDVP is adequate for most endeavors of spectral analysis and simulations, but additional treatment is needed due to the high degeneracy of methane states. Thus, we introduced the first-order Rosenkranz parameter, $Y_{n}$, to account for line mixing effects [96]. From a computational perspective, $Y_{n}$ exhibits a strong correlation with line intensity $S$. Consequently, an independent measurement of line intensity is crucial for acquiring dependable line-mixing parameters. Together with the intensity measurements by NIST and DLR, a global-fit procedure was developed based on the LMSDV model. We also converted the measured nitrogen-broadening coefficients to air-broadening coefficients using the formula $\gamma_{0}^{\text{air}}=0.79\gamma_{0}^{\mathrm{N}_{2}}+0.21\gamma_{0}^{\mathrm{O}_{2}}$, and utilizing polynomial parameters for $\gamma_{0}^{\mathrm{O}_{2}}$ as provided by Lyulin et al. [57]. The nitrogen pressure shifting coefficients were similarly scaled as $\delta_{0}^{\text{air}}=0.79\delta_{0}^{\mathrm{N}_{2}}+0.21\delta_{0}^{\mathrm{O}_{2}}$, again utilizing parameters for $\delta_{0}^{\mathrm{O}_{2}}$ as provided by Lyulin et al. [57]. Values of $Y_{n}^{\mathrm{N\textsubscript{2}}}$ for the measured CH4–N2 system are converted for CH4–air through multiplication by 0.985, as shown in Tran et al. for the $2\nu_{3}$ band [113]. The speed-dependent Voigt parameters were incorporated in HITRAN2024 wherever available.
@@
Broadening parameters ($\gamma_{0}$) were gathered from all post-year 2000 sources [3, 6, 17, 21, 22, 20, 25, 26, 27, 30, 33, 34, 35, 39, 58, 59, 62, 63, 65, 81, 79, 80, 91] that we could find, and separated by collision partners, transition symmetries, and ranking index. Unlike for other rotors, only the ranking index, $\alpha$, lets us differentiate between lines of the same manifold and symmetry, which simply relates to the energy level instead of a motion. Because of this, the parameters do not exhibit any clear pattern with $m$ for each index (see Fig. 14). Thus, separation was only used to isolate the most intense lines, which are found in $\alpha=1$.
@@
The different line profiles used in these studies impact the value of the retrieved coefficients by a small but non-negligible amount. For consistency with the parameters already present in HITRAN, we chose to focus on the Voigt profile. Separating each profile would greatly reduce the amount of data per fit; therefore, we scaled the respective non-Voigt parameters by a constant value to match the Voigt broadening. The constant scaling was calculated by taking the ratio of Voigt to non-Voigt broadening for each transition and averaging the result. This approach could only be performed in cases where transitions were fit with both non-Voigt and Voigt profiles —thus not all studies could be included. To further increase the amount of available data, the CH4–N2 pressure broadening parameters were combined with those of –air, and scaled by a factor of 0.988 taken from Pine & Gabard [79]. This approach was previously used in other studies [65] and agrees very well with other measurements (see Fig. 15). Finally, thanks to the weak branch dependence, broadening parameters for the $P$, $Q$, and $R$ branches were combined. Ideally, each branch should be separate; however, the broadening measurements in methane are too uncertain for the difference to be reliably quantified.
@@
With its different symmetry, CH3D must be treated differently, and benefits from a more rigorous separation based on quantum numbers. The semi-empirical approach developed by Dudaryonok et al. [23] seems to produce very good results in the $\nu_{6}$ band. Ignoring the wavenumber dependence, we applied their method for the broadening and temperature-dependence parameters of every new line added in Sec. 2. The wavenumber dependence of the pressure shift was reintroduced, summing the semi-empirical value to the empirical formula in HITRAN2012 [11] of $2\times 10^{-6}\times\Delta\nu$, limited to negative values. A similar approach was used where the extremely low or high values of HITRAN2020 were replaced along with the new lines, which only amount to about 70.
@@ 4 Conclusion
The parameters of more than 80,000 lines of 12CH4, 13CH4, and 12CH3D have been updated for HITRAN20204 using the new line lists provided. Transitions wavenumbers, intensities, and lower-state energy levels were taken directly from these lists to replace the values from HITRAN2020. Laboratory spectra in the pentad, octad and tetradecad regions were used by the HITRAN team to validate and choose between the overlapping data sets, while other regions relied on the validations from the other groups. We observed a great drop in the residuals of the octad and tetradecad regions spectra particularly thanks to the atm.161 list and the new data from Reed et al. [82] and Yin et al. [122]. The effects on the pentad region were limited, but still showed improvements, especially at high er energies. 45,000 lines of the various isotopologues have been added, including some lower-intensity CH3D lines that should have been in the previous update but were missing, and increasing the region covered from 12,000 to 14,000 cm-1 with experimental data. In regions that were previously empirically determined, this work allowed us to reassign some of the intensities to the correct isotopologues. Assignments of many lines have been completed between the octad and triacontad regions.
The broadening parameters of these newly added lines were replaced using Padé-approximant functions fit from multiple recent studies for 12CH4 and 13CH4, unless reliable experimental data were provided, such as those generated at the USTC. The functions were obtained from data sets separated based on their rotational symmetry and ranking index. Currently, only one function per symmetry was used, regardless of the ranking index. The pressure-induced shifts were calculated using the same formula from HITRAN2012 [11]. Semi-empirical functions from Dudaryonok et al. [23] were used for 12CH3D. The pressure shifts were scaled using the HITRAN2012 formula to account for the vibrational dependence, while the broadening parameters were taken as is. The broadening parameters from the $\nu_{3}$ band, were replaced similarly using the same Padé approximants, however, experimental values were used when available.
@@
The relational structure of the HITRAN database [45] and functionality of HAPI [54] allow one to add non-Voigt line shape parameters. For instance, in this edition, we have added the speed-dependent Voigt parameters and first-order line mixing for air-broadening in the tetradecad region. It will be important to have a consistent set of parameters in all the bands as was done for CO2 [44], N2O, and CO [43] in HITRAN2020. Of course, it is much harder for methane as there are strong symmetry and vibrational dependencies in comparison with linear molecules. Moreover, it will be important to introduce the full line mixing parameters, where possible. The accuracy of first-order line mixing deteriorates quickly at pressures approaching $1$ atm and above. Dense spectral regions such as $Q$ branches are especially affected and need full line mixing, possibly in combination with speed-dependent effects [16, 6], to be properly measured. Unlike some other parameters, these require complete knowledge of the assignments to be implemented; therefore, more theoretical work will have to be done in preparation. Currently, data from Tran et al. is available for CH4–air [112, 113] and other collision partners [111, 110], and will be considered in future updates.
In addition, in order to support the studies of planetary and brown dwarf atmospheres, it is important to provide parameters associated with broadening by “planetary” gases, including H2O [104], H2, He and CO2 [105]. Unlike many other gases for which these “planetary” broadening parameters are available, only broadening by water vapor is available for methane. In this work, we have used the scaling factors recommended in Ref. [104] to assign broadening by water vapor values to all new lines of methane. We have collected existing information on broadening by hydrogen, helium, and carbon dioxide (including Refs. [6, 110, 115, 117, 116, 29, 31, 25, 33, 118, 103, 123, 47, 18], but it is still a lot of work to process these data, which often do not agree with each other within stated uncertainties.